OpenStudy (anonymous):

what is the slope of the line 2x + 5y = -10?

OpenStudy (anonymous):

put the equation into slope-intercept form by isolating y y = mx + b m = slope

OpenStudy (anonymous):

nope, not quite correct 2x +5y = -10 5y = -2x -10 y = (-2/5)x - 2 m = -2/5